# LoRA vs Full Fine-tuning: An Illusion of Equivalence

### v1

@@ 3 Model Differences Between LoRA and Full Fine-Tuning
Figure 5: Impact of cosine similarity threshold $\epsilon$ on the number of intruder dimensions. Here, we set $k=10$ and measure the impact of $\epsilon$ on the number of intruder dimensions measured. LoRA introduces many intruder dimensions in the top 10 ranking singular vectors, while full fine-tuning does not. Top row is for RoBERTa-base.
Figure 6: Very high rank LoRA updates still have lower effective rank than full-finetuning. This holds across all matrix types and layers. (Left) Effective rank LoRA and Full. (Right) Zoomed in on only LoRA.
Full fine-tuning updates have a higher effective rank than LoRA updates, even when LoRA is performed with a full-rank matrix. Another way we can examine differences between LoRA and full fine-tuning is to calculate the effective rank (Roy & Vetterli, 2007) of the change made to the weights during fine-tuning. As shown in Fig. 6, when we calculate this we observe that the effective rank of full fine-tuning solutions have a significantly higher effective rank than solutions learned by LoRA, even when LoRA has high rank. Even at high adapter ranks and with rank stabilization, we find across layers that the effective rank of LoRA updates is less than half that of full fine-tuning and a quarter of the adapter rank. For example, with the high rank of $r=768$ for RoBERTa, LoRA updates have an average effective rank of 300. This suggests that LoRA is under utilizing its full capacity $r$, and may help explain observed gaps between LoRA and full fine-tuning on challenging tasks like coding (Biderman et al., 2024; Zhuo et al., 2024).
Intruder dimensions are distributed across both high and low singular values. We examine the extent to which intruder dimensions exist throughout the entire weight matrix and how they are distributed. To do this, we hold $\epsilon$ fixed and measure the number of intruder dimensions while varying the proportion of the fine-tuned singular vectors that we examine. We report these results in Fig. 11(a). Here, we can see that LoRA consistently has more intruder dimensions than full fine-tuning, regardless of what fraction of the singular values we examine. The only caveat to this is that, for some datasets, full fine-tuning passes LoRA with rank 1 when examining the last 20% of the fine-tuned singular vectors. This is likely due to the limited expressivity of rank 1 updates and is interesting because it suggests that in these cases, full fine-tuning may be changing lower-ranking singular vectors more than LoRA.
Intruder dimensions ncrease in magnitude and change in direction as fine-tuning progresses. To further understand how a particular intruder dimension is introduced during fine-tuning with LoRA, we measure the maximum cosine similarity between the top individual fine-tuned singular vectors and all the pre-trained singular vectors across many intermediate steps in the fine-tuning process, as seen in Fig. 7 (left). In parallel, we track changes in their associated singular values as seen in Fig. 7 (right). As is evident from the graphs, intruder dimensions appear to gradually increase their “rank” (on the left) as their singular value is increased (on the right) while simultaneously changing in direction too as training progresses.
@@
Figure 7: Evolution of the intruder dimension with training iterations. (Left) Intruder dimensions, and their rank, in a LoRA fine-tuned weight matrix during fine-tuning. (Right) Their associated singular values. This clearly shows that across training steps, the impact of the intruder dimension, as determined by its singular value, increases.
Scaling $\alpha$ with the rank of the LoRA update reduces the number of intruder dimensions alongside increasing the effective ranks of the matrices. Following (Biderman et al., 2024), we set $\alpha=2r$. However, we ran additional experiments with a fixed $\alpha=8$, as in most early work on LoRA. This has the effect of scaling down the LoRA contribution as rank increases. We report these results in Appendix 18. For both settings of $\alpha$, models obtained equivalent performance on the target task (see Table 1). With fixed $\alpha$, however, all ranks of LoRA—even very large ones—exhibit intruder dimensions. Furthermore, when we measure the effective rank of these models, they have a much smaller effective rank than when $\alpha=2r$. This suggests that with constant $\alpha$, LoRA converges to a low rank solution. This provides additional evidence that $\alpha=2r$ improves the solution of high ranks of LoRA(Kalajdzievski, 2023; Biderman et al., 2024): it leads to a reduction in intruder dimensions and an increase in the effective rank of solutions when LoRA’s rank is higher.
The total number of intruder dimensions increases proportionally to the size of the fine-tuning dataset. Using the training described in Appendix A, we fine-tuned models on data subsets of varying sizes. We trained RoBERTa-base on MNLI using LoRA with rank 1 and 8 (cases where we originally saw intruder dimensions). We then again measure number of intruder dimensions along with the impact of $\epsilon$ and $k$, and report our results in Appendix 12. For $r=8$, as we train on more data, more intruder dimensions are introduced. Interestingly, however, LoRA with rank 1 appears to converge to similar amounts of intruder dimensions, regardless of the dataset size. This may be because of the limited expressivity of models with $r=1$.
Conjecture: Intruder dimensions, as high-ranking singular vectors, contribute significantly to the norm and stability of the parameter matrix. In contrast to pre-trained singular vectors that are learned from large pre-training corpora, LoRA introduces intruder dimensions learned solely from the smaller dataset of the fine-tuning task, which overpower the pre-trained vectors, as seen in the experiments so far. On the other hand, full fine-tuning, while adapting just as effectively to the fine-tuning task, retains the spectral properties of the pre-trained model effectively. From this, we conjecture that the presence of intruder dimensions in LoRA models has a detrimental effect on the model’s performance outside the fine-tuning task distribution and this effect is less pronounced in full fine-tuned models. We investigate this conjecture in the next section.
4 Behavioral Differences Between LoRA and Full Fine-Tuning
We have identified structural differences in the solutions of LoRA and full fine-tuning. Here, we investigate whether LoRA and full fine-tuning produce measurable differences in fine-tuned model behavior. While we have already seen that they perform nearly identically on their in-distribution test set, we evaluate whether these behavioral similarities hold under other distributions.
Figure 8: Continual Learning performance of RoBERTa for full fine-tuning and LoRA. We sequentially train on six tasks, in order from left to right. Horizontal dotted line indicates baseline pre-trained performance. Vertical solid line indicates when a specific dataset is fine-tuned on. Gray region represents performance before the model has been trained on that task. We are interested in the differences in accuracies of these methods both right after training (at the vertical black line) and later (in the white region). We see that low ranks of LoRA forget previously learned tasks more.
@@ 4 Behavioral Differences Between LoRA and Full Fine-Tuning
At lower ranks, LoRA adapts less robustly during continual learning by forgetting more of the previous tasks. We train RoBERTa sequentially on multiple tasks and measure how much performance changes as new tasks are learned. We use the same training recipe and datasets as before, but now instead fine-tuning in a continual learning environment with the following dataset order: MNLI, QQP, SST-2, SIQA, Winogrande, FEVER. After training on a certain dataset in the sequence, we merge the LoRA weights into the model and reinitialize them before training on the next task so that they are unimpacted by the previous tasks. After training on a specific task, we test on all tasks by, for each task, separately retraining its classification head before testing on its test set. This enables us to examine how well the model performs on these tasks while not actually changing the model itself.
Results are shown in Fig. 8. While LoRA matches the performance of full fine-tuning initially, smaller ranks of LoRA consistently show greater degradation of performance during continual learning. In particular, we note that for the first three datasets trained on, performance of LoRA when $r=1$ drops below the pre-trained baseline. As we increase the rank of LoRA, we can see that this forgetting behavior decreases and more closely resembles full fine-tuning and even forgets less on MNLI after the completion of continual learning. Biderman et al. (2024) describe a family of tasks and training procedures under which LoRA forgets less than full fine-tuning, these results show that the complete picture is nuanced: while in some cases LoRA appears to forget less, for some tasks—and some ranks—LoRA may in fact forget more.
Figure 9: RoBERTa’s performance on its pre-training data distribution after fine-tuning on a particular task. We measure pseudo loss as described by Salazar et al. (2020). All the models for a specific task were trained to equivalent performance. We see a U-shaped curve that identifies the best rank for learning the downstream task while forgetting the pre-training distribution the least.
For LoRA models fine-tuned to equivalent test accuracy, we see a U-shaped curve that identifies the optimal rank for fitting to the downstream task while forgetting the pre-training distribution the least. We measure the shift in performance that our fine-tuned models, trained to equivalent test accuracy, have on their pre-training data distribution. While we cannot directly measure a true perplexity on encoder-only style models because they are not auto-regressive language models, we can still measure their pseudolikelihood on pre-training data as described in Salazar et al. (2020). We measure “pseudo-loss” for all our fine-tuned RoBERTa models across the four datasets that RoBERTa used during pre-training (openwebtext (Gokaslan & Cohen, 2019), cc_news (Hamborg et al., 2017), stories (Trinh & Le, 2019), and bookcorpus (Zhu et al., 2015)), and weigh them proportionally to their contribution as described by Liu et al. (2019). We report our measured pseudo-loss scores in Fig. 9. In it, we can see a U-shaped trend between full fine-tuning and LoRA with $r=768$. Since all models achieve equivalent test accuracy, this U-shaped trend across a specific dataset identifies the optimal ranks for fitting to a down stream task distribution, and seems to point to $r=64$ as the choice that minimizes forgetting of the pre-training distribution. We can see that both a rank very low rank ($r=1$) and a very high rank ($r=768$) lead to greater forgetting on the pre-training distribution relative to full fine-tuning, while for $r=64$ we see less. That is: models fine-tuned with LoRA when $r=1$ suffer from intruder dimensions and appear to have more forgetting than $r=64$ which had no intruder dimensions. However, models fine-tuned with LoRA when $r=768$ also exhibit worse forgetting, suggesting that due to their overparameterization they are overfitting to the adaptation task. With $r=8$ and $r=64$, which are more frequently used, forget less than full fine-tuning, while ranks on either extreme forget more than full fine-tuning.
Setting $\alpha$ properly significantly impacts model performance. We continue our case study of setting $\alpha=8$ instead of $\alpha=2r$ as described in earlier sections. We repeat continual learning and pre-training forgetting experiments with fixed (rather than rank-scaled) $\alpha$, and report them in Appendix 16 & 17. LoRA models, regardless of rank, forget much more of both the pre-training distribution (MNLI, QQP, FEVER) and previously learned tasks during continual learning, highlighted by the fact that all LoRA ranks drop below baseline performance for the first two datasets. These results resemble earlier findings for $r=1$, and further suggests that when $\alpha=8$ instead of $\alpha=2r$, solutions converge to a solution with more intruder dimensions structurally and one that is behaviorally similar to the low-rank LoRA setting.
5 Why Do Intruder Dimensions Exist?

### v2

@@ 3 Structural Differences
Figure 4: LoRA has intruder dimensions, whereas full fine-tuning does not. Here, we set $k=10$ and measure the impact of $\epsilon$ on the number of intruder dimensions measured. LoRA introduces many intruder dimensions in the top 10 ranked singular vectors, while full fine-tuning does not. Numbers are reported are the sums across the entire model.
Our Method. To calculate the number of intruder dimensions in a specific weight matrix, we use Algorithm. 1. In it, we first compute the SVD of both the pre-trained and resulting LoRA and full fine-tuned weights. Then, for each of the top $k$ highest-ranking singular vectors, we measure its maximum cosine similarity with all of the pre-trained singular vectors. If this maximum cosine similarity is less than some threshold $\epsilon$, we classify this singular vector as an intruder dimension. Note that both $k$, the number of fine-tuned singular vectors to examine, and $\epsilon$, the cosine similarity threshold, are hyperparameters; we verify the robustness of our findings for a wide range of $\epsilon$ and $k$ values in Fig. 4 and Fig. 15 respectively. To determine the number of intruder dimensions in a specific model, we run this algorithm for each weight matrix in the model and sum the total.
@@
Figure 5: Evolution of an intruder dimension across training steps. (Left) Intruder dimensions, and their rank, in a LoRA fine-tuned weight matrix during fine-tuning. (Middle) Their associated singular values, which shows that the singular value associated with the intruder dimension increases. (Right) Test accuracy across training steps.
LoRA fine-tuned models contain high-ranking intruder dimensions while fully fine-tuned models do not. To characterize the differences in fine-tuning methods, we first evaluate the differences in the total number of intruder dimensions in the top 10 highest-ranking singular vectors ($k=10$). We repeat this procedure for a the range of $\epsilon$ values, our cosine similarity threshold. The results are presented in Fig. 4. For LLaMA2-7B, we find that models trained with LoRA contain intruder dimensions for ranks at least as high as $r\leq 256$. For RoBERTa, we consistently observe intruder dimensions for rank $r\leq 16$, even for low values of $\epsilon$. Interestingly, we observe that fully fine-tuned models, for all model sizes, almost never contain intruder dimensions in their top 10 singular vectors, even for epsilon values of about 0.6 to 0.9. This means that full fine-tuning makes smaller changes to the same set of high contribution pre-trained singular vectors, rather than introducing new singular vectors like LoRA. Importantly, the number of intruder dimensions appears to drop as rank increases past a certain threshold, suggesting that the low-rank nature, as well as the update rule of LoRA, induces them to occur. This is underscored by the $r=2048$ case of LLaMA2-7B fine-tuned on math (Fig 4(b)), which does not have intruder dimensions and instead has a very similar curve to full fine-tuning. As rank increases past a threshold and LoRA begins to resemble a high rank update, intruder dimensions begin to disappear.
LoRA variants have intruder dimensions. We examine 4 other LoRA variants to ensure that our findings to not just apply to vanilla LoRA. We examine AdaLoRA (Zhang et al., 2023), LoRA+ (Hayou et al., 2024), PiSSA (Meng et al., 2024), and VeRA (Kopiczko et al., 2024). In all of these cases, we find intruder dimensions with similar characteristics to vanilla LoRA (see Fig. 23). This shows our findings hold to other variants. For more discussion about these methods, see Appendix O.
Intruder dimensions are distributed across both high and low singular values. We examine the extent to which intruder dimensions exist throughout the entire weight matrix and how they are distributed. To do this, we hold $\epsilon$ fixed and measure the number of intruder dimensions while varying the proportion of the fine-tuned singular vectors that we examine (Appendix I, Fig. 15). Here, we can see that LoRA consistently has more intruder dimensions than full fine-tuning, regardless of what fraction of the singular values we examine. See Appendix I for more discussion.
Intruder dimensions increase in magnitude and change in direction as fine-tuning progresses. To further understand how a particular intruder dimension is introduced during fine-tuning with LoRA, we measure the maximum cosine similarity between the top individual fine-tuned singular vectors and all the pre-trained singular vectors across many intermediate steps in the fine-tuning process, as seen in Fig. 5 (left). In parallel, we track changes in their associated singular values as seen in Fig. 5 (middle). As is evident from the graphs, intruder dimensions appear to gradually increase their “rank" (left) as their singular value is increased (middle) while simultaneously changing in direction too as training progresses.
Additional empirical observations. 1. We find that the random seed used by LoRA to initialize its adapters plays no role in the resulting structure (see Appendix N). 2. We observe that the total number of intruder dimensions increases linearly with respect to the size of the fine-tuning dataset up to a certain point before saturating (Appendix J). 3. We study the effective rank of these fine-tuning updates (Appendix H). However, we find that this measure does not suffice to explain the behavioral differences we observe in LoRA and full fine-tuning. Also, it is important to note that even if it had, its global nature would prevent the precise examinations we conduct in future sections, like in Fig. 8.
Experimental and theoretical justification for why intruder dimensions occur. It is important to note that intruder dimensions are an empirical observation of LoRA. We find that a variety of factors play a role in the introduction of intruder dimensions. In the next section, we present results that suggest that learning rate and LoRA’s $\alpha$ contribute to intruder dimensions. In the appendix, we present findings that suggest that tuning the B matrix only leads to fewer intruder dimensions (10), and demonstrate how the addition of orthogonal vectors to the pre-trained weight matrix models the introduction of intruder dimensions well (A.2).
4 Model Differences: Forgetting and Out-of-Distribution Generalization
(a) LLaMA2-7B.
(b) RoBERTa-base.
Figure 6: LoRA forgets less, even with same fit to fine-tuning task. For LLaMA2-7B, forgetting is measured on unrelated tasks, as described in Biderman et al. (2024). For RoBERTa, Pseudo loss on a sample of its pre-training distribution measured as described by Salazar et al. (2020). In both, LoRA forgets less than full fine-tuning.
LoRA forgets less. We measure the change in out of distribution performance (forgetting) induced by fine-tuning. For LLaMA2-7B, we follow Biderman et al. (2024) and measure forgetting as the average score on Hellaswag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), Arc-Challenge (Clark et al., 2018). For RoBERTa-base, we measure its “pseudo-loss”, which is analogous to language modelling loss for encoder-only models, as described by Salazar et al. (2020) on a sample of its pre-training dataset (as described by Liu et al. (2019)). Going forward, we refer to these values as “forgetting” and report them in Fig. 6. We observe that across all tasks, full fine-tuning forgets more of its pre-training language modeling ability in comparison to LoRA. Importantly, all our RoBERTa-base models fine-tune to equivalent accuracy on the downstream task (Table 2). This extends the finding that LoRA learns less Biderman et al. (2024) to the case where LoRA and full fine-tuning have equal fit, showing that LoRA forgetting less is not simply a function of it underfitting the fine-tuning task in comparison to full fine-tuning (like in Biderman et al. (2024)), but rather a characteristic of LoRA itself.
LoRA $\alpha$ impacts generalization and intruder dimensions. For our experiments, we use the commonly used $\alpha=2r$ (Biderman et al., 2024) as well as $\alpha=8$ (Hu et al., 2021). For both settings of $\alpha$, models obtain equivalent performance on the target task (Tables 1 & 2). However, when $\alpha=8$, all ranks of LoRA—even very large ones—exhibit intruder dimensions (Fig. 19(a)), have a much smaller effective rank than when $\alpha=2r$ (Appendix H), and have much worse generalization (more forgetting, Fig. 20). Models trained with $\alpha=2r$ have fewer intruder dimensions and generalize better. This provides additional evidence highlighting the importance of using $\alpha=2r$ (Kalajdzievski, 2023; Biderman et al., 2024), particularly for higher ranks of LoRA.
An increase in intruder dimensions leads to an increase in forgetting. We do a learning rate sweep for RoBERTa-base with LoRA $r=8$ on MNLI to observe its impact. Across epochs, we measure the number of intruder dimensions, test accuracy, and forgetting (pre-training loss). We report the results of these models and our baseline full fine-tuning model in Fig. 7. We observe that across and within training runs, as the number of intruder dimensions increase, forgetting (meaning worse generalization) also increases. Test accuracy has no such relation. Separately, we see that for large learning rates with many intruder dimensions, LoRA models forget more than full fine-tuning. This shows that while LoRA in general does forget less than full fine-tuning, it is not a guarantee.
Intruder dimensions strongly correlate with forgetting. When we measure the Spearman correlation between the number of intruder dimensions with forgetting in Fig. 7, we find an extremely strong fit($\rho=0.971$, p-value $\ll 0.001$). When measuring the same for our LLaMA2-7B models across training epochs (Fig. 11), we still find a strong and still statistically significant relationship ($\rho=0.59$, p-value $=0.0006$). In contrast, when measuring the correlation between intruder dimensions and test accuracy, we find no statistically significant relationship: for RoBERTa, we measure $\rho=-0.3381$ p-value $=0.218$. For LLaMA2-7B, we measure $\rho=-0.3178$ p-value $=0.0869$. See Appendix F for more information. These results suggest that intruder dimensions are clearly linked with forgetting but are not necessary for performance. We examine this claim and whether this relationship is causal in the next section.
Figure 7: As training progresses, models with growing amount of intruder dimensions continue to forget more, despite non-increasing test performance. We also measure a strong correlation($\rho$ = 0.971, p-value $\ll 0.001$) between number of intruder dimensions and pre-training pseudo loss. Bigger learning rates lead to more intruder dimensions and forgetting.
5 Intruder Dimensions Cause Forgetting
Previously, we observed that the number of intruder dimensions correlates strongly with forgetting. Do intruder dimensions cause this forgetting?
Scaling the magnitude of intruder dimensions. To test if intruder dimensions cause increased forgetting, we must intervene on intruder dimensions and see the impact. We do this by finding the highest ranked (by singular value) intruder dimension in each weight matrix and scale its contribution such that the new weight matrix is $W=W_{0}+\Delta W+(\lambda-1)u_{i}\sigma_{i}v^{T}_{i}$, where $i$ is the index of the top intruder dimension ($\lambda=0$ is removal and $\lambda=1$ is no change). We sweep $\lambda^{\prime}s$ between 0 and 1, scale the intruder dimensions, and measure test accuracy and pre-training loss. For a comparison baseline, we select the neighbor of the intruder dimension to separately scale. See Fig. 8 for results (and Figs. 1213 for full results.)
(a) LLaMA2-7B
on code with r=16.
(b) LLaMA2-7B
on math with r=16.
(c) RoBERTa-base
on MNLI with r=8.
(d) RoBERTa-base
on QQP with r=8.
Figure 8: Scaling down intruder dimensions in fine-tuned models reduces forgetting but not performance. We scale the top intruder dimension in each matrix such that $W=W_{0}+\Delta W+(\lambda-1)u_{i}\sigma_{i}v^{T}_{i}$. Lines represent forgetting (red) and learning (green). Dotted lines represent pre-trained baselines. Axis Labels: Green: Test Accuracy (%). Red: Pre-training Loss (Forgetting).
Scaling down intruder dimensions reduces forgetting. In Fig. 8, we show that when we scale down the top intruder dimension of each weight matrix, we measure a significant reduction in forgetting (pre-training loss) while incurring a minimal drop in test accuracy. For all examples in Fig. 8, we observe that when using $\lambda=0.7$ or $0.9$, there is almost no impact on fine-tuning performance, while there is a large percentage drop in forgetting (See Tables 3&4). In one example for LLaMA2-7B fine-tuned on MetaMath with LoRA $r=256$, we observe that scaling the top intruder dimension in each matrix with $\lambda=0.3$ leads to a 0.1% drop in test accuracy and a 33.3% drop in the forgetting induced by fine-tuning. In another for RoBERTa-base fine-tuned on QQP, using $\lambda=0.7$ leads to equivalent in test accuracy and a 33.2% reduction in the forgetting induced by fine-tuning. In certain scenarios, we even see test accuracy improve along with a drop in forgetting. If we instead increase their contribution ($\lambda>1$), we observe more forgetting. Across the board, scaling down intruder dimensions seems to have little impact on test accuracy but a major impact on forgetting. This pattern is exclusive to intruder dimensions (Fig. 12): if we instead intervene on pre-trained singular vectors that have a similar singular value to the intruder dimension (ensuring similar contributions to matrix) and scale their magnitude down, we see that forgetting goes up. These results indicate that the forgetting observed in LoRA is caused by intruder dimensions interfering with the pre-trained language modeling capabilities. Moreover, the scale (singular value) of these intruder dimensions is not essential for the fine-tuning task performance. See Appendix G for further discussion.
@@ 5 Intruder Dimensions Cause Forgetting
Continual Learning Setup. To examine a practical example of how accumulating intruder dimensions, which cause forgetting, may impact performance, we study continual learning, since it requires learning and remembering across a range of tasks. To do this, we train RoBERTa sequentially on multiple tasks and measure performance as new tasks are learned. We use the same training recipe and datasets as before but now sequentially in the following dataset order: MNLI, QQP, SST-2, SIQA, WinoGrande, FEVER. After training on a certain dataset in the sequence, we merge the LoRA weights into the model and reinitialize the LoRA adapter before training on the next task. After training on a specific task, we test on all tasks by, for each task, separately retraining its classification head before testing on its test set. Results are shown in Fig. 9(a).
Accumulating intruder dimensions hurts LoRA models during continual learning. In Fig. 9(a), initially both LoRA and full fine-tuning train to equal performance (MNLI), which is consistent with our previous observations. However, we observe that all ranks of LoRA degrade much more rapidly than full fine-tuning. Low ranks of LoRA, which have the most intruder dimensions, degrade the most. We attribute this divergence from earlier results—where LoRA appeared to forget less—to the accumulation of intruder dimensions during continual learning, which drive forgetting. To show this, we visualize how intruders are added across tasks in Fig. 9(b). Here, we see that each task adds its own intruder dimensions leading to a large amount of intruder dimensions upon the completion of the six task continual learning experiment. In contrast, in Fig. 9(c), we see that full fine-tuning retains the pre-trained structure well justifying why full fine-tuning forgets less during continual learning.
(a) Continual learning results.
(b) Similarity matrices for LoRA r=8 during continual learning.
(c) Similarity matrices for full fine-tuning during continual learning.
Figure 9: Full fine-tuning is better than LoRA at continual learning because of accumulating intruder dimensions. When sequentially training on six tasks, full fine-tuning retains performance better than LoRA. in Fig. 9(a), horizontal dotted line indicates baseline pre-trained performance. Vertical solid line indicates when a specific dataset is fine-tuned on. Gray region represents performance before the model has been trained on that task. See Appendix L, Fig. 17 for more. In Figs. 9(b)&9(c), we see that LoRA accumulates intruder dimensions across tasks and contributes to its degrading performance, whereas full fine-tuning does not.
Implications and prescriptions in fine-tuning. These findings suggest several implications and prescriptions during LoRA fine-tuning. We have shown that intruder dimensions drive forgetting, and therefore should be avoided when possible. Interestingly, this presents a data free model evaluation method to examine which model is most overfit to the fine-tuning task (forgotten the most): given two equally performing models on downstream test sets, you should select the one with fewer intruder dimensions. Intruder dimensions appear to be a necessary part of fine-tuning, but they can be mitigated. Further, these results show the danger of using LoRA during continual learning and justifies using many different adapters without combining them, like advocated in Sheng et al. (2024).
6 Conclusion
This paper describes the finding that LoRA and full fine-tuning update different parts of the parameter space resulting in distinct spectral properties: LoRA often introduces intruder dimensions—high-ranking singular vectors dissimilar to those in pre-trained weights. These structural differences persist across a series of ablations. Next, we find that models with fewer intruder dimensions exhibit better out-of-distribution generalization and forget less of the pretraining distribution. Last, we show that intruder dimensions cause increased forgetting: We show that reducing the magnitude of high ranking intruder dimensions leads to minimal changes in test performance but a large drop in pretraining loss. We show that this is particularly relevant during continual training: even though LoRA forgets less than full fine-tuning after training on one task, sequentially training leads to an accumulation of intruder dimensions that causes more forgetting than full-finetuning.
Acknowledgements
We would like to thank Jacob Portes and Dan Biderman for corresponding with us and releasing their LLaMA-2 7B checkpoints for us to use. This enabled us to study a more comprehensive range of models. We would also like to thank Leshem Chosen, Lucas Hennigen, Han Guo, Vighnesh Subramaniam, Valerio Pepe, and the entire Language & Intelligence lab for their helpful feedback on this work. This research was supported in part by the National Science Foundation under grant IIS-2238240.
Appendix A Why do intruder dimensions exist & can we alleviate them?
Here, we discuss possible causes of intruder dimensions.
A.1 What do intruder dimensions do?
Conjecture: Intruder dimensions, as high-ranking singular vectors, contribute significantly to the norm and stability of the parameter matrix. In contrast to pre-trained singular vectors that are learned from large pre-training corpora, LoRA introduces intruder dimensions learned solely from the smaller dataset of the fine-tuning task, which overpower the pre-trained vectors, as seen in the experiments so far. This suggests that these intruder dimensions are very task specific. On the other hand, full fine-tuning, while adapting just as effectively to the fine-tuning task, retains the spectral properties of the pre-trained model effectively. Our experiments that scale down intruder dimensions provide evidence for the claim that intruder dimensions are specialized to the fine-tuning task, since we observe that increasing the norm of an intruder dimension (using $\lambda>1$) leads to little change in adaptation task performance but leads to a significant increase in forgetting (pre-training loss).
A.2 Adding a random vector